OpenStudy (hang254):

The lengths of the sides of a rectangle become 4 times greater than the original measurements. Which statement is true about the rectangle? Its area becomes 4 times greater, and perimeter becomes 20 times greater. Its area becomes 16 times greater, and perimeter becomes 4 times greater. Its area becomes 4 times greater, and perimeter becomes 16 times greater. Its area becomes 20 times greater, and perimeter becomes 4 times greater

OpenStudy (ajprincess):

Let the original length and width be l and w. then the area=l*w perimeter=2(l+w) Nw the lengths are made 4 times greater. So new area becomes 4l*4w=16lw new perimeter =2(4l+4w) =2(l+w)*4 So Its area becomes 16 times greater, and perimeter becomes 4 times greater is the correct option. Is that clear?

OpenStudy (hang254):

Thank You, that was very helpful!

OpenStudy (ajprincess):

yw.